If fn converges to f and FN is bounded on a set S prove that {fn} is uniformly bounded on S
1
Expert's answer
2020-05-18T18:47:21-0400

We need to assume that "f_n\to f" uniformly on "S". Hence, "\exist\, N\in\mathbb N \, \forall m,n\geqslant N \colon |f_n(x)-f_m(x)|<1". Also "\exist\, C_1,C_2,\dots>0 \colon |f_n(x)|\leqslant C_n,\, x\in S,\,n\in\mathbb N." So, for "n\geqslant N,\, x\in S" we have "|f_n(x)|\leqslant|f_n(x)-f_N(x)|+|f_N(x)|<1+C_N". Taking "C=\max\{C_1,\dots,C_N\}+1" we have "\forall n\in\mathbb N\colon |f_n(x)|\leqslant C" because "\max\{C_1,\dots,C_N\}" bounds "|f_n|" for "n<N".